OpenStudy (anonymous):

find the equation of the line passing threough the points ( -1,7) and (2,-8). Write the equation in slope intercept form. I know I have to use M =y2 - y1 / x2 - x 1 Using that I got -8 - 7 / 2 - -1 That gave me y = -15 x = 3 -15/3 = -5 Now what?

OpenStudy (anonymous):

Use the standard form for the line : y=mx+b You got m=-5---------------------->y=-5x+b. then you can pick the point (-1,7) or (2-8) for the line passes. Plug the value of x, y to find b then the write the line with the value of m and b

OpenStudy (anonymous):

which one is m and which one is B?

OpenStudy (anonymous):

-5 = m and -5 = y?

OpenStudy (anonymous):

m=-5 Let pick the point (-1,7)------>7=(-5)(-1)+b 7=5+b b=7-5=2 Line: y=-5x+2

OpenStudy (anonymous):

Thanks! :D its 3:17 am and im learning math